OpenStudy (anonymous):

solve using square roots. x2 - 10 = 26

OpenStudy (anonymous):

add 10 on both sides

OpenStudy (anonymous):

i think its \[\huge x^2-10x=26\]

OpenStudy (anonymous):

\[x ^{2}-10+10=26+10\]

OpenStudy (anonymous):

your solving for square roots.

OpenStudy (anonymous):

\[x ^{2}=36\]

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

@alijahe I think x=-6 should also be one of the answers